What is proven by the result obtained with the reaction of silver nitrate with the water in the beaker? With I2/KI? What is the meaning of semi-permeability? What is permeable and what is not? What determines permeability?
1
Expert's answer
2020-11-09T14:01:19-0500

4AgNO3 + 2H2O → 4Ag + O2 + 4HNO3

in this case, no hydrolysis occurs. Silver cation does not bind hydroxide anion

2AgNO3 + 2HOH = Ag2O + 2HNO3

2Ag (+) + HOH = Ag2O + 2H (+)

silver hydroxide AgOH at the moment of formation decomposes into oxide and water: 2AgOH = Ag2O + H2O

For this reason, AgNO3 salt is stored in a dry place in dark packaging.


As a result of the interaction of potassium iodide with silver nitrate, a yellow crystalline precipitate of silver iodide and potassium nitrate is formed. The molecular equation of the reaction is:

KI+AgNO3=AgI+KNO3

Let us write down the ionic equations, taking into account that silver iodide does not decompose into ions, i.e. does not dissociate.

Ag (+) + NO3 (-) + K (+) + I (-) = AgI+K (+) + NO3 (-)

Ag (+) + I (-) = AgI

The reaction with silver nitrate is a qualitative reaction for chloride, bromide and iodide ions.

I2 + AgNO3 → AgI + INO3


Permeability is a passage or diffusion of a gas, vapour, liquid, or solid through a material without physically or chemically affecting it.

Semi-permeable means that it allows some substances/molecules to pass through, but not others


For example, light is permeable in air, an impenetrable fabric for water and where